The second side of a triangle is 1 centimeter more than three times the first side. The third side is 2 centimeters longer than the second side. If the perimeter is 46 centimeters, find the lengh of each side of the triangle
Answer by Nate(3500)

The second side of a triangle is 1 centimeter more than three times the first side. The third side is 2 centimeters longer than the second side. If the perimeter is 46 centimeters, find the lengh of each side of the triangle.
First Side: s
Second Side: 3s + 1
Third Side: 3s + 1 + 2 = 3s + 3
(s) + (3s + 1) + (3s + 3) = 46
s + 3s + 1 + 3s + 3 = 46
7s = 42
s = 6
First Side: s = 6
Second Side: 3s + 1 = 3(6) + 1 = 19
Third Side: 3s + 1 + 2 = 3s + 3 = 3(6) + 3 = 21
6 x 19 x 21